OpenStudy (anonymous):

log_6(x)+log_6(x+1)=1

OpenStudy (unklerhaukus):

Combine the logs on the left like this \[\log_b(N)+\log_b(M)=\log_b(N\times M)\]

OpenStudy (anonymous):

Yeah I got log_6(x)(x+1)=1 so far

OpenStudy (unklerhaukus):

Now use the definition of a log \[\log_bz=y\quad\iff\quad z=b^y\]

OpenStudy (unklerhaukus):

can you do this part?

OpenStudy (kc_kennylau):

That means do the 6 to the power both sides

OpenStudy (unklerhaukus):

yes,

OpenStudy (kc_kennylau):

@MikeB55 your turn now :)

OpenStudy (shamim):

log_6(x^2+x)=log_6 6 @MikeB55

OpenStudy (unklerhaukus):

you'll get a quadratic which has two solutions, only one of these solutions solves the original equation

OpenStudy (kc_kennylau):

Looks like Mike's not here, let's wait for his return :)

OpenStudy (anonymous):

Hey sorry everyone

OpenStudy (anonymous):

I'm at x^2+x=6

OpenStudy (anonymous):

So x=2

OpenStudy (kc_kennylau):

Yep

OpenStudy (anonymous):

Thanks everyone

OpenStudy (kc_kennylau):

no problem

OpenStudy (unklerhaukus):

that's right!, did you also get the number that solves the quadratic but not the log equation? do you know why this solution doesn't work?

OpenStudy (kc_kennylau):

I assume he already knows it :)

OpenStudy (anonymous):

Yeah I got it!

OpenStudy (shamim):

why -3 is not accepted @Unkle

OpenStudy (kc_kennylau):

Because log of negative is undefined

OpenStudy (anonymous):

Negative numbers can't be used to solve a log equation

OpenStudy (shamim):

ok

OpenStudy (unklerhaukus):

exactly